OpenStudy (kenialove):

The product of three consecutive whole numbers is 990. What is the sum of the three whole numbers ? A. 30 B. 33 C. 63 D. 96

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do you know what "consecutive" means in this question ?

OpenStudy (misty1212):

maybe?

OpenStudy (kenialove):

No not really @misty1212

OpenStudy (misty1212):

one after another like say \(10,11,12\) are consecutive so lets find the three numbers ok?

OpenStudy (kenialove):

Okay

OpenStudy (kenialove):

How doe ?

OpenStudy (misty1212):

pick three numbers and multiply them together, see if you get 990

OpenStudy (misty1212):

since it ends in a zero, one of them is probably 10

OpenStudy (kenialove):

Any numbers ?

OpenStudy (misty1212):

so we can try \[10\times 11\times 12\] and see if it works no, not any numbers, CONSECUTIVE numbers

OpenStudy (kenialove):

Yea that's wat I mean any pair of consecutive numbers

OpenStudy (misty1212):

\[10\times 11\times 12=1230\] too but, not right have to go smaller

OpenStudy (misty1212):

it is 3 numbers, not a pair

OpenStudy (kenialove):

U know wat I mean

OpenStudy (kenialove):

9 x 10 x 11 = 990

OpenStudy (misty1212):

ok got it on the second try good now add them up!

OpenStudy (misty1212):

\[9+10+11=?\]]

OpenStudy (kenialove):

30

OpenStudy (kenialove):

Oohhh I get it thanks so much @misty1212

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]